In a ground state atom of Br how many electrons total will have the quantum number l=2?
1
Expert's answer
2020-07-07T14:22:28-0400

"l=2"

Magnetic quantum number from "(-l \to +l)" ":(m_1)=-2,-1,0,1,2" "(" Total "=5)"

For magnetic quantum number,there are two spin quantum number that are "(+\frac{1}{2},-\frac{1}{2})"

So total number of electrons for"\ l=2 \implies 5\times 2=10"